The concentrations of gaseous species in the react
Answer #1

P 0.05 atm

Kp = 3.9*10^-2

and

KP = PNH3*PH2S

since the other reactant is solid, it is not included

then

K = 3.9*10^-2

then

KP = PNH3*PH2S = 3.9*10^-2

in equilibrium

PNH3= PH2S

x^2 = 3.9*10^-2

x = sqrt(3.9*10^-2) = 0.1974841

PT = 0.1974841+0.1974841 =0.3949682

this is slightly favoured toward products

then, this is NOT in equilbirium